### v1

# Revolutionizing Command Interface: Maximizing Control Efficiency in INO ICAL Experiment with UDP Protocol

Abstract
Efficient command interface is a critical requirement for experiments employing a large number of front-end DAQ modules and control servers. In the context of the INO ICAL (India-based Neutrino Observatory Iron Calorimeter) experiment, this involves 28,800 Resistive Plate Chamber (RPC), charged particle detectors. The acquisition and control of these detectors are facilitated through Front End data acquisition modules known as RPC-DAQs. These modules consists of Ethernet interfaces for data and command connectivity to a server. Each module acts as a network node with a unique IP address. The collective group of hundreds of modules is controlled by a common server over a Local Area Network (LAN). UDP (User Datagram Protocol) is the most commonly used networking protocol which supports Multicast as well as Unicast, can be easily adapted to INO ICAL Experiment. A server can send commands to group of DAQs or any particular DAQ. But UDP may have the problem of packet loss and reliability. To mitigate these issues, this paper suggests a simpler approach that modifies the UDP protocol by implementing a handshaking scheme and checksum, similar to those found in more reliable protocols like TCP. The proposed solution optimizes the use of UDP as a reliable command interface in the INO ICAL experiment, ensuring seamless data acquisition and control. Also, this paper shows the performance study of the custom hybrid UDP Command Interface in the prototype ICAL experiment called Mini Iron Calorimeter (Mini ICAL) which houses 20 units of RPCs and electronics. This work not only addresses the challenges of the INO ICAL experiment but also underscores the adaptability and robustness of the proposed protocol for usage in mini-ICAL and beyond.
1 Introduction
Data acquisition and control are the crucial components of the readout electronics in the particle physics experiments. Within the context of the INO-ICAL (India-Based Neutrino Observatory Iron Calorimeter) experiment, a substantial number of sensors are employed known as RPCs (Resistive Plate Chambers). Each of these RPCs is interfaced with a digital front end known as RPC-DAQ [1] (RPC Data Acquisition Module) as shown in Figure 1. The ICAL experiment uses a vast array of 28,800 individual RPCs with RPC-DAQ Front End, each serving as a detector unit.
Figure 1: RPC-DAQ position in INO-ICAL experiment.
These DAQ modules with wired Ethernet interface [2] are positioned near the detectors they serve as shown in Figure 1. Each of these RPC-DAQs are configured with a unique IP address. This web of DAQ modules along with the interconnected detectors, collectively forms what is commonly referred to as the sensor network nodes. The primary objective of ICAL experiment is to track charged particles over a large area, so it is important to ensure the continuous and synchronized operation of these detectors. Central command servers are developed to oversee the group performance of RPC-DAQs modules. In principle these servers act as the control and configuration nodes for the entire ICAL.
Network communication with efficient protocol design is crucial in ICAL experiment. Standard TCP/IP networking protocols are readily available, but the demands of the experiment, factors such as timing precision, data reliability, and group management, require a careful and tailored approach. Due to this, it is clear that customizing standard protocols is the ideal approach to meet the needs of the ICAL experiment.
2 Selection of Protocol and Customization
In choosing the right protocol for the ICAL project, which comprises 28,800 RPCs, several critical factors must be considered. The connection less and light weight nature of User Datagram Protocol (UDP) is well suited for ICAL experiment. It supports both Multicast and broadcast, which are essential for group management. But UDP lacks data integrity while data transfer. On the other side, the Transmission Control Protocol (TCP) supports reliable data transfer and error-checking mechanism. But TCP is not preferred when it comes to multicast and broadcast functionalities as shown in Figure 2, making it less suited for efficient group management in the ICAL project.
Figure 2: Types of UDP over TCP Protocol.
By combining the reliability features of TCP with the group managing features of TCP a hybrid protocol was designed called Hybrid Protocol based Command Interface (HPCI). HPCI improves the standard connection less UDP protocol with handshaking and error detecting capabilities of TCP, supporting a balanced and efficient protocol that supports unique demands of ICAL project. HPCI is used as a command-and-control server for not only RPC-DAQ modules but also to other subsystems like Trigger and Data Concentrator to synchronise the data acquisition in ICAL experiment.
3 HPCI Model Overview
In the HPCI model, multiple DAQs are interconnected with a central command server as shown in Figure 3. Users have the capability to issue global commands from the server, which can be distributed to all the connected DAQs through Multicast grouping. Each DAQ establishes a socket in Multicast mode, effectively registering to a shared Multicast Group IP address, often denoted as (e.g., 239.0.0.1), utilizing protocols like IGMP/ICMP.
Figure 3: Prototype HPCI Model with 5 RPC-DAQs.
Apart from Multicast the HPCI model supports transmission of commands to individual DAQs through Unicast mode. For example, initiating a "RUN START" command can be sent in Multicast mode, making sure that the same command packet reaches all DAQs simultaneously. In situations where a particular DAQ shows irregular behavior during a run, users can switch to Unicast mode to request the status of that specific DAQ. To support this feature, each DAQ must have two open sockets—one used to handle Multicast packets and the other handles Unicast packets. Acknowledgments of both types of commands are sent to the server through the Unicast socket.
The pilot HPCI model infrastructure was developed to handle 100 RPC-DAQs simultaneously. Ethernet switches are used at both Level 1 and Level 2, to handle Gigabit data transfer and communication efficiency. All the RPC-DAQs are configured with IPv4 addresses to support standard network protocols. Each DAQ has the feature of remotely programmable network configuration, allowing users to modify network configurations when needed. Network Configuration like IP Address, MAC Address, GATEWAY and SUBNET are permanently stored in the flash memory of each DAQ . During the boot sequence, the FPGA processor reads network information from a dedicated location of its Flash memory. This helps in configuring correct network information every time while DAQ is booting. Dedicated commands are available with user to change the network data in the flash memory.
4 Command Flow between Front-end and Back-end
HPCI has been implemented as a central command server to oversee the control of all the DAQs. The HPCI command execution follows sequence of steps in both Front-end and Back-end as illustrated in Figure 4. A dedicated UDP server is setup to issue user selected command from the back-end. Receiving these commands the corresponding DAQ carries out series of tasks in the front-end and sends back an acknowledgement. CRC error checking is done for both command and acknowledgement packets. In case of negative or no acknowledgment the command is retransmitted in Unicast mode and the DAQ database is updated with only accessible DAQ list.
For Example: To execute a status command called "ISDAQUP", at first a local database is used to read the installed DAQ list. Then the command server sends this status command in Multicast mode to all the DAQs. Also the UDP socket on the back-end is always kept in listening mode to capture acknowledgments from all connected DAQs. When receiving "ISDAQUP" command the DAQ sends an acknowledgment packet which consists of DAQ-ID. After receiving acknowledgments, the command server updates the local DAQ database. The "ISDAQUP" command will be retransmitted in Unicast mode to all DAQs that have not responded within the specified timeout period. During the run start sequence, series of commands are issued to all the DAQ nodes listed in the updated database.
Figure 4: Command Transfer between back-end and Frontend.
The command flow process starts with a group command followed by sequence of Unicast commands. The payload size varies depending on the command type, as depicted in Figure 7 and listed in Table 3. Some of the UDP command packets don’t have payload because they are used only to retrieve DAQ information. The UDP command packet size varies from minimum 18 bytes to maximum 100 bytes. Each packet has a field called "Seq no" which is incremented every time when a retransmission occurs after no acknowledgement. This is used to keep track of network health of all the RPC-DAQs.
5 Implementation and Testing of the HPCI
To validate and assess the functionality of the HPCI model, a prototype network system was setup on test bench. This prototype comprised of 5 RPC-DAQs and a central server, as illustrated in Figure 3. The network was organized using a Multicast IP address, 239.0.0.1. This test setup serves as a realistic network architecture to quantify various parameters. Two distinct test methods are devised a Busy Model, characterized by the intense acquisition of TCP based event data from all 5 RPC-DAQs, and a Non-Busy Model as shown in Table 1 and Table 2, wherein the network load was primarily attributed to command transmissions only.
Table 1: Cycle time measurement Non-Busy Model.
Command Type DAQ Command Average Cycle time
Processing Time ($\mu$s) of a command ($\mu$s)
Command A (18 Bytes) 174 300
Command B (26 Bytes) 405 537
Command C (100 Bytes) 615 813
Table 2: Cycle time measurement Busy Model.
Command Type Network Load (Mbps) Average Cycle time of a command ($\mu$s)
Command C (100 Bytes) 6kHz * 600Bytes = 27 854
Command C (100 Bytes) 10kHz * 600Bytes = 45 864
It was seen that, the network response time exhibited a marginal increase during periods of high network activity. It’s worth noting that the test results presented herein were conducted under one-to-one correspondence between the DAQs and the server. In real operational conditions, the network is expected to experience substantial congestion due to the collection of TCP event data from all DAQs. This results in response times exceeding one millisecond. These test results will help in adjusting timeout settings in final experiment.
5.1 Command Processing at the Back-end
Managing commands in the back-end server of the HPCI model involves communicating with large number of DAQs simultaneously. Also the back-end command server is crucial for receiving acknowledgments from these DAQs and keeping a detailed database to monitor their health status. Also the back-end server controls the retransmision of non-acknowledged commands.
Command processing at the back-end are executed sequentially including steps like reading database, setting up UDP server, selecting payload and adding CRC word. For example, when the user needs to find the active status of all the DAQs, a task accomplished by invoking the ’ISDAQUP’ Command. At the start the run control process consults the pre-loaded DAQ database, extracting information and formulating an expected DAQ list (D). the back-end server creates a dedicated UDP socket to transmit the command in Multicast mode. The Acknowledgment reception process then captures acknowledgment responses from each DAQ, facilitating the creation of an active DAQ list (R). A comparative analysis of lists D and R enables the back-end to compile a roster of failed DAQs (F). The retransmission process takes charge of each DAQ on the F list, reissuing the same command for a specified number of iterations (n). Finally, a conclusive DAQ database is constructed, based on the presently available DAQs. This database equips users with the information to initiate subsequent commands. This sequential operation shown in Figure 6 allows users to make informed decisions based on real-time DAQ status information.
Figure 5: Command Processing in Back-end.
Figure 6: Command execution in Front-end.
5.2 Command Processing at the Front-end DAQ
[…]
5.3 Critical Command Processing and Error Detection
The command interface plays an important role in executing detector specific tasks within the DAQ. Every command packet has a "Command Word" field as illustrated in Figure 7 and Table 3. Depending on the command word, the RPC-DAQ performs a specific task that alters crucial aspects of data acquisition. While some commands are fundamental for health monitoring, in scenarios where these tasks are exceptionally critical, cannot be duplicated, and involve highly sensitive parameters (where loading incorrect parameters could lead to system damage), the HPCI incorporates command log services and CRC checksum to address these challenges effectively.
Figure 7: Command and Acknowledgement Packet.
Table 3: Command Data Structure.
Command word Description
Start Marker Command : 0xDDDD , Acknowledgement : 0xEEEE
Reserved Reserved for later use in ICAL
RPC-DAQ ID DAQ number configured in each RPC-DAQ
Data Type 0xFF00/BB00+$0$/00AA (Forward/Backward+ NoACK/Ack )
Command Word Command Identification word for various command
Sequence No Increments when the command is resent by the Command server
Data Size Size of the payload Data 0…N
Data 0 to N Actual message (variable in size depends on the command)
Checksum CRC-16 Calculated for the whole packet before transmitting on either side
The command log services maintains a comprehensive record of past commands specific to each DAQ. These logs are used for analysing DAQ information required at each time a when command is dispatched by the HPCI. These logs receive continuous updates triggered by both negative and positive acknowledgments received from each DAQ. To enhance the reliability of command interface CRC (Cyclic Redundancy Check) checksum and retransmission procedures have been introduced.
CRC-16, is a well-established 16-bit checksum technique, used on the command processing paths of both Front-end and Back-end. It uses the 0x8005 polynomial for checksum word generation and error detection. Apart from CRC-16, various other checksum techniques such as Fletcher, longitudinal redundancy checks, and simple integer addition were studied. But, CRC-16 has surpassed error detection efficiency of other checksum schemes. If error is detected during command or acknowledgment processing, the respective device is retransmitted with the same command.
6 HPCI in Mini-ICAL
Mini-ICAL [4], located at the IICHEP Transit Campus in Madurai, Tamil Nadu, India, serves as a prototype for the proposed ICAL. Mini-ICAL is a scaled down version, approximately 1/600th the size of ICAL. It consists of 11 layers of iron plates with each layer housing two RPCs and their respective electronics in each slot. One of the objectives of Mini-ICAL is to address engineering challenges particularly in the context of data acquisition and detector performance in the stray magnetic field of ICAL.
Figure 8: Run Control Software of Mini-ICAL Experiment.
RPC-DAQ is strategically positioned to acquire and process data from the detectors. These modules receive power, Ethernet connectivity, global clock signals and triggers from the Back-end  [5]. To facilitate communication and networking a 24-port 10Gbit Ethernet Switch operates as an L1 Network Switch, establishing a local area network (LAN). Additionally, Mini-ICAL integrates various subsystems and servers into this LAN, connecting them via another L2 switch. The heart of Mini-ICAL’s data acquisition system lies in its Run control software shown in Figure 8 which uses HPCI for efficient command server to the RPC-DAQs and subsystems. This software controls various components of Mini-ICAL by distributing commands and managing command logs, retransmissions and error detection. The run control software comprises two windows or tabs. One window allows for the modification of network and detector configurations, while the other controls the run start stop sequence and display of useful data. Command acknowledgments are highlighted using color coding where green denotes success, and red indicates failure. A run console is included, where the software logs and displays details of each command execution. After its first installation in 2018 the run control software is being used for day to day data taking in Mini-ICAL. Throughout this period, several key issues were identified and addressed, particularly adjustments made to acknowledgment timeouts to enhance the software’s performance.
7 Other Responsibilities of Command server in ICAL Experiment
HPCI allows the command server to configure run conditions including setting up TCP sockets for event acquisition, enabling periodic monitoring, and facilitating remote system upgrades. These processes are essential for seamless data collection and system maintenance. HPCI is instrumental in distributing epoch time to all the DAQs. This timestamp is used by the Real Time Clock (RTC) hardware logic in RPC-DAQ to precisely synchronize commands and events. This helps in accurate event time-stamping and data correlation by the event servers. The Detector parameters like high voltage (HV) and channel masking are configured and monitored through HPCI. This allows the detectors to operate with specified parameters and maintain optimal condition. HPCI allows user to configure network parameters online. These settings includes IP addresses, Multicasting addresses, port numbers, and socket mode settings (TCP/UDP). HPCI provides a centralized interface for seamless network configuration adjustments.
8 Conclusion
The HPCI approach is not a novel concept in network-based DAQs, its application within the INO ICAL project presents unique challenges due to the sheer scale of network-enabled DAQs involved. So HPCI is devised with traditional command interface approach flavored with smarter handshaking techniques. Currently, HPCI based command server is scaled and tested upto 20 RPC-DAQs in mICAL. However, the real test lies ahead as it will soon be deployed and evaluated with a significantly larger number of DAQs within the INO ICAL project. The HPCI continues to evolve and time tested in Mini-ICAL.
Acknowledgments
We sincerely thank all our present and former INO colleagues especially Puneet Kanwar Kaur, Umesh L, Anand Lokapure, Aditya Deodhar, Sagar Sonavane, Suraj Kole, Salam Thoi Thoi, Rajkumar Bharathi for technical support during the design, Also we would like to thank members of TIFR namely S.R. Joshi, Piyush Verma, Darshana Gonji, Santosh Chavan and Vishal Asgolkar who supported testing and commissioning. Also we like to thank former INO directors N.K. Mondal and V. M Datar for their continuous encouragement and guidance. .